OpenStudy (anonymous):

sqrt(x+2) +sqrt(x)=4

OpenStudy (anonymous):

x = 49/16

OpenStudy (anonymous):

Square both sides and solve for x in the linear equation

OpenStudy (anonymous):

can you show me please

OpenStudy (anonymous):

I will show the solution, one minute.

OpenStudy (anonymous):

\[\sqrt{x+2}+\sqrt{x}=4\]\[\sqrt{x}=4-\sqrt{x+2}\]\[x=(4-\sqrt{x+2})^2\] \[x=x-8\sqrt{x+2}+18\] \[\sqrt{x+2}=9/4\] x=49/16

OpenStudy (anonymous):

can explain how you got x=x-8 sqrt (x+2)+18